OpenStudy (anonymous):

Find the equation of the tangent to the curve. y= x sqrt. (3x+1) when x = 5 I just need help with differentiating!

OpenStudy (anonymous):

\[y= x \sqrt {(3x+1)}\]

OpenStudy (anonymous):

Can you do it in the form \[u = \sqrt {3x+1} ---> u' =\]\[v = x -----> v' = \] It makes it easier for me to understand... What's the diferentiation of u?

OpenStudy (callisto):

u' =(1 /2) [sqrt (3x+1)] ^(-1/2) (3) = 3/2[sqrt (3x+1)] ^(1/2)

OpenStudy (callisto):

v' = 1

sam (.sam.):

y=(copy left)(diff. right)+(copy right)(diff left) \[y=x[ (1/2)(3x+1)^{-1/2} (3) ] + (\sqrt{3x+1})\]

OpenStudy (anonymous):

Why wouldn't it be u= sqrt {3x+1} ----> u' = (1/2)(1/ sqrt 3x+1) v= x -----> v' = 1 ?

OpenStudy (anonymous):

Ok, but why times by 3?

OpenStudy (anonymous):

@order , chain rule: